3) A social worker reports that 30% of workers in a factory are below 15 years of age. Of the 120 employees surveyed, 38 said they were below 15 years old. Using ? = 0.05, interpret the p-value.

1
Expert's answer
2022-06-06T13:49:30-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.30"

"H_a:p\not=0.30"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05

," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:



"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{38/120-0.3}{\sqrt{\dfrac{0.3(1-0.3)}{120}}}=0.3984"


Using the P-value approach:

The p-value is "p=2P(Z>0.3984)= 0.690335," and since "p=0.690335>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than 0.30, at the "\alpha = 0.05" significance level.